OpenStudy (anonymous):

Find 3 consecutive odd integers such that the sum of the first and the third equals the sum of the second and 31.

OpenStudy (anonymous):

2k-1,2k+1,2k+3 now you have equasion 2k-1+2k+3=31+2k+1<=> 4k+2=2k+32 <=> 2k=30 So the consecutive numbers are 29,31,33 This method is based on equal conversion which allows you to not chech whether this is indeed valid values, however if you want to be safe you can check that 29+33 is indeed 31+31 . Hope that helps. Also if you want you can mark the integers as 2k+1,2k+3, and 2k+5 however i find my marking more gentle.

OpenStudy (anonymous):

So what's the equation ??

OpenStudy (anonymous):

i meant 29, 31 and 33 would work :)

OpenStudy (anonymous):

I have written it XD Basicly just read the first four lines, the rest is my jabbertalking since i am a bit of perfectionist in math :3

OpenStudy (anonymous):

The equasion you want is 2k-1+2k+3=31+2k+1 while 2k-1,2k+1,2k+3 are the integers you are serching for

OpenStudy (anonymous):

Thank you very much :)

OpenStudy (anonymous):

No problem bro :3

OpenStudy (anonymous):

:)